### v2

# Quantum Vacuum energy as the origin of Gravity

Abstract
[…]
I Introduction
Gravity remains the least understood of the fundamental forces, in spite of the discoveries of General Relativity (GR) and Quantum Mechanics (QM) over 100 years ago, this year being recognized as the 100-th anniversary of a complete theory of QM due to Schrödinger and Heisenberg, although the birth of QM goes back to Planck’s black body studies at finite temperature. Classical singularities at the center of black holes and at the time of the Big Bang have not yet been understood and resolved in classical GR. Many physicists expect that such singularities will be resolved by a theory of quantum gravity, but it’s fair to say that a complete theory of quantum gravity has not been forthcoming, although string theory offers some promise. For these reasons it is worthwhile to pursue new ideas on the origin or emergence of Gravity itself, especially if such ideas suggest that perhaps Gravity doesn’t need to be quantized in the first place. Such a theory should be able to explain the value of Newton’s constant $G_{N}$ itself.
In this article we explore the question “Can Quantum vacuum energy be the origin of Gravity?” [^2] This article was initially intended simply as a theoretical exercise in the form of an essay to address this question, but we found it can lead to a fruitful new perspective on Gravity more generally, and it is not without some predictions, if the basic ideas continue to hold. This perspective is based on the idea that perhaps we can learn more about the foundations of Gravity by considering the cosmological far future, in contrast to the far past where much less is understood, in particular the origin of the Big Bang and possibly inflation. The late Universe is dominated by so-called Dark Energy, thus one should first attempt to understand the full implications of this in a Universe where Dark Energy is its sole component. By assumption, we will equate Dark Energy with the quantum vacuum energy density $\rho_{\rm vac}$} computed in quantum field theory (QFT) in flat local Minkowski space, defined as
$\langle 0|T_{\mu\nu}|0\rangle={\rho_{\rm vac}}\,\eta_{\mu\nu}$ (1)
with the convention $\{\eta_{\mu\nu}\}={\rm diag}\{1,-1,-1,-1\}$ where the Vacuum $|0\rangle$ is the ground state for all quantized fields, excluding gravity. The gravitational implications of the quantum field theoretic properties of $\rho_{\rm vac}$} will be then explored.
The first observations of Gravity were of course at the surface of the Earth which is manifested as a constant acceleration $g$ for all matter regardless of mass or specific material. Newton later understood $g$ to be “environmental” rather than fundamental, namely $g=G_{N}M_{E}/R_{E}^{2}$, where $M_{E},R_{E}$ are the mass and radius of the Earth. In Newton’s universal theory of Gravity, $G_{N}$ is considered a new fundamental constant of Nature, and this view was adopted by Einstein in his formulation of General Relativity. [^3] Throughout this paper, $G_{N}=6.67\cdot 10^{-11}{\rm m^{3}}/{\rm kg\,s^{2}}$ is Newton’s constant as measured today. We keep the fundamental constants $\hbar$ and $c$ explicit in order to make certain conceptual points and to provide some numbers with units. “Natural” units based on the Planck length $\ell_{p}$ can be deceptive and are not so natural for this article. In contrast, below we will argue that $G_{N}$ is also environmental, namely
$G_{N}=\dfrac{c^{4}}{2}\dfrac{{{\rm R}_{\scriptscriptstyle\infty}}}{{{\rm E}_{\scriptscriptstyle\infty}}}=\dfrac{c^{2}}{2}\dfrac{{{\rm R}_{\scriptscriptstyle\infty}}}{{{\rm M}_{\scriptscriptstyle\infty}}}$ (2)
[…]
Based on the above remarks, what will be central to our exercise is the quantity $\rho_{\rm vac}$}. For purposes of illustration, consider a Universe which only consists of a single harmonic oscillator with ground state energy $E_{0}=\hbar\omega/2$, in addition to an imagined observer. Since classical mechanics in unaffected by shifts of the potential energy by an arbitrary constant, $E_{0}$ is not measurable in classical physics. If all that exists is the harmonic oscillator in empty space, then $E_{0}$ also cannot be measured in a single experiment based on QM. There are two known ways to measure $E_{0}$. First, one can couple the harmonic oscillator to a generic heat bath at temperature $T$. Then one can in principle infer $E_{0}$ from the high temperature behavior of the partition function or its logarithm which is the free energy. It’s important that this is based on the free energy and not its derivatives, and thus it is not invariant under shifts of the potential energy. The second manner to measure $E_{0}$ is to couple the oscillator to Gravity, since gravity originates from all the matter-energy in the Universe and is sensitive to its ground state energy. The fact that the ground state energy can be measured either by coupling to a heat bath or gravity suggests parallels between thermodynamics and gravity that have already been recognized. Deceptively simple as it may appear, to study QFT at finite temperature with periodic boundary conditions, one necessarily needs to study quantum fields in curved Euclidean space. In 1 spatial dimension the Euclidean space is a torus, and the problem of conformal theory on a torus is a completely solved problem. Like Gravity, the Vacuum $|0\rangle$ is omnipresent. The Vacuum “knows” the physical properties, such as mass, quantum numbers, and interactions of any particle excitations above it, and is perhaps the only entity that exists everywhere in the Universe and for all times. Everything else is some quantum excited state, including people, and typically unstable. Nature does not abhor a Vacuum, rather it depends on it. [^5] To paraphrase with translation a recognizable philosopher, If you look deeply enough into the Void, it eventually looks back at you. So it is reasonable to suppose that the Vacuum, in particular its ground state energy density $\rho_{\rm vac}$}, is at the origin of gravity. In other words, if the only way to measure $E_{0}$ is with gravity, then reversing the argument, perhaps quantum vacuum energy is the origin of Gravity. For the same reason that one does not necessarily have to quantize the heat bath to study finite temperature QFT, perhaps one does not have to quantize Gravity, however this deep issue will be left as an open question. This does not preclude the existence of classical gravitational waves that have been measured, as they can be viewed as classical perturbations of the Vacuum as an elastic medium. This article explores this possibility as a theoretical academic exercise to see where it leads.
This article assumes that $\rho_{\rm vac}$} is well-defined in Minkowski space, namely it is as well defined as the ground state energy of a harmonic oscillator. Early considerations of $\rho_{\rm vac}$} based on viewing a free quantized field as a collection of harmonic oscillators with energies $\hbar\omega_{\bf{k}}$, with $\omega_{\bf{k}}=\sqrt{{\bf{k}}^{2}+m^{2}}$ led to the introduction of a cut-off in $|{\bf{k}}|$. If this cut-off is taken to be based on the Planck length $\ell_{p}\equiv\sqrt{\hbar G_{N}/c^{3}}$ then $\rho_{\rm vac}$} is much too high by even 100 orders of magnitude and this is perhaps the first formulation of the “cosmological constant problem” Weinberg . This naive computation of $\rho_{\rm vac}$} depends on the cut-off and goes to infinity as the cut-off goes to infinity. This computation is essentially meaningless, as is the often quoted “worse prediction in physics by $10^{120}$ orders of magnitude”, simply because it is not a real prediction. In quantum field theory one is accustomed to dealing with ultra-violet (UV) singularities. In recent work ALrhovac1 ; ALrhovac2 we formulated a proper computation of $\rho_{\rm vac}$} which renders it finite and well-defined. Since this will be central to subsequent arguments we list its main properties.
(i)   For 2 dimensional integrable theories $\rho_{\rm vac}$} is exactly computable from the Thermodynamic Bethe Ansatz, and also from the S-matrix bootstrap, and takes the form ${\rho_{\rm vac}}=m_{1}^{2}/2\mathfrak{g}$ where $m_{1}$ is the lightest mass particle. In the works ALrhovac1 ; ALrhovac2 this formula was extended to 4 spacetime dimensions:
${\rho_{\rm vac}}=\dfrac{3}{4}\dfrac{c^{5}}{\hbar^{3}}\dfrac{m_{\rm z}^{4}}{\mathfrak{g}}$ (3)
where $m_{\rm z}$ is the physical mass of the lightest particle, which we will refer to as the “zeron”, where zero signifies the Vacuum, and $\mathfrak{g}$ is a dimensionless interaction coupling constant. Although one does not have the powerful simplifications of integrability in 4 dimensions, this formula was well motivated in ALrhovac1 ; ALrhovac2 and should not be viewed as an ad hoc assumption, although unlike in 2 dimensions it hasn’t yet been fully rigorously proven since the $1/\mathfrak{g}$ dependence indicates it is highly non-perturbative. Clearly the above expression is purely quantum mechanical as can be seen from its $\hbar$ dependence. The $1/\mathfrak{g}$ dependence can be anticipated, and is in fact necessary, since $\rho_{\rm vac}$} should diverge as $\mathfrak{g}\to 0$ due to an unavoidable logarithmic divergence for free QFT in the UV when the spatial dimension $d$ is evenALrhovac1 . [^6] This is essentially another, but refined, version of the “Cosmological Constant Problem”. The precise identification of the zeron as a real particle will not be necessary for our arguments, thus this identification will be left as an open question. However it was argued in ALrhovac2 that $m_{\rm z}$ is fixed in the deep UV above the electro-weak spontaneous symmetry breaking scale, and that Majorana neutrinos are viable candidates (see (8)). [^7] For experimental constraints on neutrino masses see for instance NeutrinoMasses .
[…]
${\cal F}=-\dfrac{\pi^{2}T^{4}}{90}\,c(r),~~~~~~{\rm with}~~c(r)=c_{\rm uv}+c_{4}\,r^{4}+\ldots$ (4)
for some constant coefficients $c_{\rm uv},c_{4}$. [^8] In 2 spacetime dimensions $c_{\rm uv}$ is the effective Virasoro central charge. Then ${\rho_{\rm vac}}=-c_{4}\pi^{2}m_{\rm z}^{4}/90$. This was used in ALrhovac1 to estimate $\rho_{\rm vac}$} for QCD based on finite temperature lattice results.
(iii)   $\rho_{\rm vac}$} can be independently computed from the form-factor bootstrap that relates the two particle form-factor to the zero particle one, which is a 1-point vacuum expectation value of the field in question ALrhovac2 . This is the reason for $m_{\rm z}$ to be the lightest mass particle, since in principle the S-matrix and form factors for other particles can be obtained by bootstrapping the lightest mass particle. Applying this to $\langle 0|T_{\mu\nu}|0\rangle$, this leads to a determination of the coupling $\mathfrak{g}$. As usual, defining $\cal T$} from the S-matrix $S=1+i{\cal T}$, then
$\lim_{s\to\infty}{\cal T}(s)=\dfrac{\mathfrak{g}}{m_{\rm z}^{2}}\dfrac{1}{s},$ (5)
where $s=(p_{1}+p_{2})^{2}$ is the Mandelstam variable for 2-particle scattering.
In two spacetime dimensions, the two above definitions of $\rho_{\rm vac}$} were shown to agree exactly for a variety of integrable quantum field theories ALrhovac1 ; ALrhovac2 . As previously mentioned, in one spatial dimension it takes the form ${\rho_{\rm vac}}=c^{3}m_{1}^{2}/2\hbar\mathfrak{g}$ where $m_{1}$ is the lightest particle and $\mathfrak{g}$ can be explicitly computed. Various consistency checks on this formula were made, including that it must vanish for supersymmetric theories and also theories with a fractional supersymmetries where the hamiltonian is in the center of the universal enveloping algebra of the conserved charges. (See ALrhovac1 and references therein.) The coupling $\mathfrak{g}$ can be positive or negative, and $\rho_{\rm vac}$} can even oscillate in sign for a given model, for instance for the sine-Gordon model where it oscillates from positive to negative, where it is zero at the fractional supersymmetric points. In 4 spacetime dimensions it was not rigorously shown that equations (4), (5) lead to the same formula for $\rho_{\rm vac}$}, thus (3) should still be viewed as well-motivated but still conjectural. For the remainder of the article we simply assume the well-motivated formula (3).
Some independent support for the result (3) comes from Swampland ideas Festina1 ; Festina2 . By studying a charged particle of mass $m$ in the presence of a black hole, it was argued that
${\rho_{\rm vac}}<\dfrac{m^{4}}{2e^{2}}$ (6)
where here $\mathfrak{g}$ is identified with the quantum electrodynamic (QED) fine structure constant $\alpha=e^{2}/\hbar c$. Actual QED will not play a role in this article since we assume the zeron is electrically neutral, however, as we will see, it will be instructive to compare couplings and renormalization group parameters with QED, just to check if they are reasonable. Our result (3) is stronger than (6) since it is not an inequality, but it is still consistent with (6).
As previously stated, in this article we do not propose what the hypothetical zeron particle actually is, nor even that it is a detectable particle, however we can constrain it’s mass $m_{\rm z}$ from the cosmologically measured value of $\rho_{\rm vac}$}. Taking it to be ${\rho_{\rm vac}}=\Omega_{\Lambda}\rho_{\rm crit}$, where $\Omega_{\Lambda}\approx 0.68$ one finds Planck
${\rho_{\rm vac}}=5.2\times 10^{-10}\,\dfrac{{\rm Joule}}{{\rm m}^{3}}.$ (7)
Using this value, then (3) implies
$m_{\rm z}=0.0024\,\mathfrak{g}^{1/4}\,{\rm eV}.$ (8)
For $\mathfrak{g}={\cal O}(1)$, it is interesting to note that $m_{\rm z}$ is on the order of proposed neutrino masses NeutrinoMasses . Below we will propose additional constraints on the coupling $\mathfrak{g}$.
It is instructive to compare the perspective on Gravity developed in this article with some proposed theories of quantum gravity, since this clearly delineates some important distinctions. We take “quantum gravity” to mean a theory where the space-time metric $g_{\mu\nu}(x)$ itself is quantized which implies the existence of spin 2 gravitons. Although the definition of the Planck length $\ell_{p}$, defined from $G_{N}$ as $G_{N}=c^{3}\,\ell_{p}^{2}/\hbar$, is a priori just based on dimensional analysis, one expects that $\ell_{p}$ is a relevant scale when considering graviton-graviton scattering. The latter does not play a role in this article, which makes one question whether $\ell_{p}$ is a fundamental constant, however this issue will not be completely settled in the present cosmological context. [^9] Based on this article, units based on $\hbar$, $c$ and $m_{\rm z}$ are more natural. A significant result of string theory is that it provides a consistent prescription for the computation of graviton scattering that is finite StringTheory . Thus gravity “emerges” from string theory, in the sense that Newton’s constant $G_{N}$ is determined by matching graviton scattering with low energy gravity. This leads to a formula for $G_{N}$:
$G_{N}\sim c_{Y}\,g_{s}^{2}\,\,\dfrac{c^{3}\ell_{s}^{2}}{\hbar}$ (9)
[…]
With the above introductory remarks, let us summarize the remainder of this article. The two primary assumptions we make are the formula (3) for $\rho_{\rm vac}$} which was well-motivated by the works ALrhovac1 ; ALrhovac2 , and the renormalization group flow for $\mathfrak{g}$ in equation (35) below. In the next section we formulate a gravitational version of the electromagnetic Casimir effect, and provide an argument for how gravity can arise from $\rho_{\rm vac}$}. In particular we show how Einstein’s field equations emerge, in this context in the form of Friedmann’s equations. The formula (2) leads to new interpretation of the Gibbons-Hawking entropy of de Sitter space, wherein the quantum informational “bits” are identified as quantized massless particles at the horizon with wavelength $\lambda=2\pi{{\rm R}_{\scriptscriptstyle\infty}}$. The analysis in Section II is independent of the form $\rho_{\rm vac}$} in (3), and we turn to understanding the implications of (3) in Section III. There we assume $\mathfrak{g}$ is a marginal coupling with a renormalization group (RG) flow, which as we will argue is well-motivated, and show that this leads to an effective, induced RG flow for Newton’s constant $G_{N}$ as a function of an energy scale $\mu$. [^11] We previously published some preliminary work on the possible implications of the RG flow of $\mathfrak{g}$ for cosmology in LeClairUniverse , however the present article supersedes it. As is standard in Cosmology, we tie this energy scale $\mu$ to the redshift $z$ and obtain equation (43). Solving the resulting equations for the scale factor $a(t)$ results in a significant modification of de Sitter space that is essentially an inverted gaussian and at no time is $a(t)=0$, which signifies no classical type of geometric singularity since for all times $a(t)>a_{\rm min}$. The solution to the scale factor satisfies an interesting symmetry between the far past and far future, namely $a(t)=a(-t+2t_{\rm min})$ where $a(t_{\rm min})=a_{\rm min}$. There is no Big Bang in our model since the solutions in past and future match at the time $t_{\rm min}$. Extending our solution to the far past, the time evolution of the Universe is more analogous to the swing of a pendulum. Our interpretation of this is that the age of the Universe is actually infinite since the solution to $a(t)$ extends to the far past without ever reaching the singularity at $a=0$. In Section IV we add matter and radiation, and due to the RG flow of $G_{N}$ in (43) there is again no $a=0$ singularity and the symmetry between the past and future continues to hold. This implies very small logarithmic corrections to the $\Lambda{\rm CDM}$ model if $\widehat{b}$ is small. We apply these ideas to the so-called Hubble tension, wherein the discrepancy of measurements of Hubble constant $H_{0}$ is proposed to arise from the fact that the two contradictory measurements rely on data that refers to different epochs where Newton’s constant is effectively different since the energy scales involved are different.
II Emergence of Einstein’s equations from a Gravitational Casimir effect
In this section we formulate a gravitational version of the electromagnetic Casimir effect. Without assuming Einstein’s field equations, we present heuristic arguments for how they emerge in this context. These arguments lead to the formula (2) for Newton’s constant $G_{N}$ expressed in terms of length and energy scales in the Universe. This doesn’t rely on the explicit formula (3) for $\rho_{\rm vac}$}, where its implications will be studied in the next section.
Let us first review the electromagnetic Casimir effect, which is well understood theoretically and has been measured Casimir . This Casimir effect requires two uncharged, parallel conducting plates in vacuum separated by a distance $L$. The energy density of quantized photons between the plates $\rho_{\rm rad}$ leads to a measurable force between the plates $F=\partial_{L}E$ where $E=\rho_{\rm rad}\cdot V$ where $V$ is the volume between the plates. It is important to point out that the theoretical predictions, which are confirmed in the measurements, do not measure the vacuum energy density $\rho_{\rm vac}$} but rather the energy density of free photons to leading order, whose pressure is $p=w\rho_{\rm rad}$, where $w=1/3$, whereas genuine vacuum energy density has w=1w=-1. In fact, since the free theory of photons is conformally invariant, it’s vacuum energy density ${\rho_{\rm vac}}=0$.
II.1 Formulating a gravitational Casimir effect without plates
For gravity, there is no well-formulated analog of the electromagnetic Casimir effect, since hypothetical gravitons cannot be constrained by material boundaries like plates. In any case, if any experiment could be imagined that overcomes these obvious difficulties, its effects would be much too small to be measurable. This led us to formulate a gravitational Casimir effect without material boundaries, namely without plates, that is sensitive to $\rho_{\rm vac}$}, where the latter is the vacuum energy density of all quantum matter and radiation excluding gravity. Throughout this paper we assume ${\rho_{\rm vac}}>0$ such that we are dealing with de Sitter space. We make three additional assumptions: (i) $\rho_{\rm vac}$} is well-defined and finite in flat Minkowski space for the entirety of quantum fields for matter and radiation without gravity, defined in (1). This quantity $\rho_{\rm vac}$} is in principle computable without gravity, as explained in the Introduction. Secondly, (ii) The Vacuum has no time dependence originating from Einstein’s equations, namely $\partial_{t}{\rho_{\rm vac}}=0$. However as we will see in the next section, tying the RG energy scale to the redshift $z$ leads to an additional induced time dependence. Third, (iii) There is no additional and independent free parameter $\Lambda$ in Einstein’s equations corresponding to a classical cosmological constant. This is consistent with the fact that $\rho_{\rm vac}$} is not invariant under arbitrary shifts of the potential energy since it is in principle computable from the logarithm of the partition function and not its derivatives.
We carry out this construction in arbitrary spatial dimension $d$ in order to make certain points. Consider first a relativistic QFT in $d=1$ spatial dimension, from which one can compute $\rho_{\rm vac}$}. There are an infinite number of integrable QFT’s where one can calculate $\rho_{\rm vac}$} exactly; see for instance the examples in ALrhovac1 ; ALrhovac2 . Let the spatial dimension be a finite size segment of length $2R$. Since total vacuum energy is $E=2R\,{\rho_{\rm vac}}$, there is a force $F=dE/dR=2{\rho_{\rm vac}}$ that can in principle be measured by moving the “walls”, in this case simply the endpoints of the segment. However if the endpoints do not consist of matter being held in place, the effect of this force is expansion or contraction of the size of the segment, depending on the sign of $\rho_{\rm vac}$}. This naturally leads us to introduce geometry, namely a metric $g_{\mu\nu}$:
$ds^{2}=-g_{\mu\nu}(x)dx^{\mu}dx^{\nu}=-c^{2}dt^{2}+a(t)^{2}dx^{2}.$ (10)
The dimensionless scale factor can be thought of as $a(t)=R(t)/R_{0}$ where $R_{0}$ is the scale factor today, such that at the present time $t_{0}$, $a(t_{0})=1$. In order to make sense of Newton’s 2nd law, one needs to introduce an inertial mass $M$. One also needs to introduce a length scale $R$ such that $R\ddot{a}$ is an acceleration, where as usual over-dots refer to time derivatives. Equating the force $F$ with $2{\rho_{\rm vac}}$, and requiring both sides scale the same way with the scale factor $a$, one obtains
$MR\,\ddot{a}=2a\,{\rho_{\rm vac}}~~~\Longrightarrow~~\dfrac{\ddot{a}}{a}=\dfrac{2{\rho_{\rm vac}}}{MR}.$ (11)
We next require $\partial_{t}{\rho_{\rm vac}}=0$. One needs another equation to enforce this, which is $\ddot{a}/a=(\dot{a}/a)^{2}$:
$\left(\dfrac{\dot{a}}{a}\right)^{2}=\dfrac{2{\rho_{\rm vac}}}{MR}.$ (12)
One can easily check that one time derivative of (12) combined with (11) implies $\partial_{t}{\rho_{\rm vac}}=0$.
Let us extend the above arguments to $d$ spatial dimensions, where now $R$ is a radius. Define $V_{d}(R)$ to be the volume of a sphere in $d$ dimensions. Then one has
$MR\,\ddot{a}=\dfrac{a}{d}\,\dfrac{\partial V_{d}(R)}{\partial R}{\rho_{\rm vac}},~~~~~V_{d}(R)=\dfrac{2\pi^{d/2}R^{d}}{d\,\Gamma(d/2)}.$ (13)
The factor of $1/d$ on the right hand side of the first equation is due to the vectorial nature of the force. To summarize, we have two equations after imposing $\partial_{t}{\rho_{\rm vac}}=0$:
$\dfrac{\ddot{a}}{a}=\dfrac{R^{d-2}}{M}\left(\dfrac{2\pi^{d/2}}{d\,\Gamma(d/2)}\right){\rho_{\rm vac}}$ (14)
$\left(\dfrac{\dot{a}}{a}\right)^{2}=\dfrac{R^{d-2}}{M}\left(\dfrac{2\pi^{d/2}}{d\,\Gamma(d/2)}\right){\rho_{\rm vac}}.$ (15)
Since $\rho_{\rm vac}$} is independent of time, $H\equiv\dot{a}/a$ is a constant. If one identifies $H=c/R$, then the second equation just implies
${\rho_{\rm vac}}\cdot V_{d}(R)=Mc^{2},$ (16)
such that $Mc^{2}$ can be interpreted as the total vacuum energy in a sphere of radius $R$.
Thus far, Newton’s constant $G_{N}$ has played no role, since we have not assumed Newtonian gravity nor Einstein’s general relativity. $G_{N}$ can be identified by comparing with Einstein’s field equations for this particular metric. We adopt standard conventions for the Einstein-Hilbert action in $d$ dimensions:
${\cal S}=\dfrac{1}{16\pi G_{d}}\int d^{d+1}x\,\sqrt{-g}\left({\cal R}+{\cal L}_{\rm matter}\right).$ (17)
The metric is as in (10), with with $dx^{2}$ replaced with $d\vec{x}\cdot d\vec{x}$. There is no a priori reason to include a curvature term proportional the standard $k\in\{-1,0,1\}$ in the spatial metric based on the above formulation of the Casimir effect. In other words $k=0$ is natural for our formulation, since, from our perspective, gravity originates from $\rho_{\rm vac}$} in Minkowski space, and the latter clearly has $k=0$. This leads to the field equations
${\cal R}_{\mu\nu}-\tfrac{1}{2}g_{\mu\nu}\,{\cal R}=\dfrac{8\pi G_{d}}{c^{d+1}}\,T_{\mu\nu}.$ (18)
If one carefully keeps track of the $d$-dependence, this leads to the following Friedmann equations [^12] The $d$ dependence of this formula should be a well-known result. One can find it for instance in Gibbons .
$\dfrac{\ddot{a}}{a}=-\dfrac{8\pi G_{d}}{d(d-1)c^{2}}\left[(d-2)\rho+d\,p\right]$ (19)
$\left(\dfrac{\dot{a}}{a}\right)^{2}=\dfrac{16\pi G_{d}}{d(d-1)c^{2}}\,\rho.$ (20)
Specializing to vacuum energy with $T_{\mu\nu}=\langle 0|T_{\mu\nu}|0\rangle$ in (1), where $p=-{\rho_{\rm vac}}$, one obtains the equations (14) with the identification
$G_{d}=c^{2}\dfrac{R^{d-2}}{M}\left(\dfrac{(d-1)\pi^{(d-2)/2}}{8\Gamma(d/2)}\right).$ (21)
[…]
It remains to specify $R$ and $M$ in the above equations. From our formulation of the gravitational Casimir effect, it’s clear that $M$ and $R$ refer to large scale properties of the Universe, and this requires some input from the solution $a(t)$. Recall that so far we have only considered vacuum energy density in $T_{\mu\nu}$, thus we are dealing with de Sitter space where $H\equiv\dot{a}/a$ is a constant by the second Friedmann equation (20). The natural choice is then
$R=R_{H}=\dfrac{c}{H},$ (22)
which is referred as the Hubble radius. With the identification of $G_{d}$ in (21), the equation (20) is equivalent to
${\rho_{\rm vac}}V_{d}(R_{H})=Mc^{2}.$ (23)
In de Sitter space, $R_{H}$ has several equivalent meanings which are all well-known and we briefly review. Here
$a(t)=e^{H(t-t_{0})}.$ (24)
$R_{H}$ equals the event horizon, which marks the boundary beyond which an observer at a given time can never receive signals from events due to the accelerating expansion. For a co-moving observer in de Sitter space, the proper distance is given by
$d_{\rm event}=a(t)\int_{t}^{\infty}\dfrac{c\,dt^{\prime}}{a(t^{\prime})}=R_{H}.$ (25)
The Hubble volume $\tfrac{4}{3}\pi R_{H}^{3}$ represents the volume that will remain causally connected to an observer as $t\to\infty$. In other words, it is the maximal volume of space that can influence an observer’s world line. $R_{H}$ also represents the distance at which objects recede in the expansion at the speed of light, which perhaps explains why the formula (2) is the correct relation for the mass $M_{\bullet}$ and radius $R_{\bullet}$ of a black hole if one identifies $M_{\bullet}={{\rm M}_{\scriptscriptstyle\infty}},R_{\bullet}={{\rm R}_{\scriptscriptstyle\infty}}$. Based on the above discussion we identify $R=R_{H}\equiv{{\rm R}_{\scriptscriptstyle\infty}}$, and $M={{\rm M}_{\scriptscriptstyle\infty}}$ defined through (23) where the subscripts refer to the far future. In 3 spatial dimensions, this gives the formula in the Introduction:
$G_{N}=\dfrac{c^{2}}{2}\dfrac{{{\rm R}_{\scriptscriptstyle\infty}}}{{{\rm M}_{\scriptscriptstyle\infty}}}.$ (26)
Let us for the moment express (26) as $G_{N}=3c^{2}H^{2}/8\pi{\rho_{\rm vac}}$. If one relies only on experimental data in modern cosmology, then the above expression (26) for $G_{N}$ is approximately correct with some assumptions, however it should be realized that at this stage this is a tautology since analysis of the cosmological data depends on $G_{N}$. First of all, we have not yet included matter and radiation as excitations over the Vacuum, where $H$ is no longer a constant, and we will return to this below. Let us simply point out that if one takes the experimental value (7) and identifies $H=\sqrt{\Omega_{\Lambda}}H_{0}$ where $H_{0}$ is Hubble constant as measured by the CMB (52) and $\Omega_{\Lambda}=0.68$, then the above expression evaluates to $G_{N}=6.72\times 10^{-11}$ ${\rm m}^{3}/{\rm kg}\,{\rm s}^{2}$. Let us note that the same computation using $H_{0;{\rm SN}}$ in (52) yields a higher value for $G_{N}$. This will be addressed below in connection with the “Hubble tension” by taking into account the explicit expression (3) for $\rho_{\rm vac}$}.
II.2 An interpretation of the Gibbons-Hawking entropy of de Sitter space
There are some important results in semi-classical gravity where the gravitational field is not quantized, namely they do not rely on the existence of gravitons, but the result still depends on the Planck length $\ell_{p}$, such as the Bekenstein-Hawking entropy and temperature for black holes Bekenstein ; Hawking . In this present context, the analog is the Gibbons-Hawking entropy for de Sitter space GibbonsHawking . Since these results depend on $G_{N}$, the expression (2) for $G_{N}$ leads to an interesting new interpretation of this entropy.
[…]
$S=k_{B}N_{\rm bits}.$ (27)
The Gibbons-Hawking entropy was thus proposed to be
$S_{GH}\equiv\dfrac{k_{B}}{4}\dfrac{A}{\ell_{p}^{2}}=k_{B}\left(\dfrac{\pi c^{3}}{\hbar}\right)\dfrac{{{\rm R}^{2}_{\scriptscriptstyle\infty}}}{G_{N}}.$ (28)
where $A=4\pi{{\rm R}^{2}_{\scriptscriptstyle\infty}}$. Here each bit is commonly interpreted as a pixelation of the area $A$ into quantized units of area $4\ell_{p}^{2}$, suggesting that spacetime itself is somehow quantized into these minimal area units at the horizon. Such a quantization of spacetime itself remains a somewhat vague notion, and difficult to make sense of mathematically.
[…]
$S=k_{B}\,\dfrac{{{\rm E}_{\scriptscriptstyle\infty}}}{E_{\lambda}},~~~~~~{\rm with}~E_{\lambda}=\dfrac{\hbar c}{\lambda},~~\lambda=2\pi{{\rm R}_{\scriptscriptstyle\infty}},$ (29)
where as above ${{\rm E}_{\scriptscriptstyle\infty}}={{\rm M}_{\scriptscriptstyle\infty}}c^{2}$. Here, the bits are quantized energies rather than hypothetical quantized patches of spacetime. Remarkably, the above formula is identical to the Gibbons-Hawking entropy, including the overall $1/4$, if one identifies $G_{N}$ as above (2):
$S=k_{B}\dfrac{{{\rm M}_{\scriptscriptstyle\infty}}c^{2}}{E_{\lambda}}=k_{B}\left({{\rm M}_{\scriptscriptstyle\infty}}c^{2}\right)\left(\dfrac{2\pi{{\rm R}_{\scriptscriptstyle\infty}}}{\hbar c}\right)=\dfrac{k_{B}}{4}\dfrac{A}{\ell_{p}^{2}},$ (30)
where in the above equation $\ell_{p}=\sqrt{\hbar G_{N}/c^{3}}$ and $G_{N}$ is given by (2). The analog of the Hawking temperature $T\equiv{{\rm T}_{\scriptscriptstyle\infty}}$ as usual follows from $d{{\rm E}_{\scriptscriptstyle\infty}}={{\rm T}_{\scriptscriptstyle\infty}}dS$ which gives the simple result
$k_{B}{{\rm T}_{\scriptscriptstyle\infty}}=E_{\lambda}.$ (31)
[…]
Let us point out that the above interpretation of the Gibbons-Hawking entropy also applies to the Bekenstein-Hawking entropy of a black hole. The reason is that (2) is the correct relation between the mass $M_{\bullet}$ of the black hole and its radius $R_{\bullet}$ if one makes the replacement ${{\rm M}_{\scriptscriptstyle\infty}},{{\rm R}_{\scriptscriptstyle\infty}}\to M_{\bullet},R_{\bullet}$:
$G_{N}=\dfrac{c^{2}}{2}\dfrac{R_{\bullet}}{M_{\bullet}}.$ (32)
From this perspective, $G_{N}$ is a universal constant which is the same for every black hole in the Universe. Repeating the above arguments for the Bekenstein-Hawking entropy,
$S_{BH}=\dfrac{k_{B}}{4}\dfrac{A}{\ell_{p}^{2}}=k_{B}\left(\dfrac{\pi c^{3}}{\hbar}\right)\dfrac{R_{\bullet}^{2}}{G_{N}}=k_{B}\,\dfrac{E_{\bullet}}{E_{\lambda}},~~~~~~{\rm with}~E_{\bullet}=M_{\bullet}c^{2}~~{\rm and}~~E_{\lambda}=\dfrac{\hbar c}{2\pi R_{\bullet}},$ (33)
where we have used (32) in the second equation. The Hawking temperature $T_{H}$ of the black hole is again given by (31), namely $k_{B}T_{H}=E_{\lambda}$.
III Induced renormalization group flow for Newton’s constant
III.1 Effective RG flow for Newton’s constant based on the RG flow for $\rho_{\rm vac}$}.
The results of the last section apply whatever $\rho_{\rm vac}$} is. In this section we use the expression (3) to obtain further results. For simplicity of subsequent discourse let us rewrite equations (19),(20) explicitly for $d=3$:
$\dfrac{\ddot{a}}{a}=\left(\dfrac{\dot{a}}{a}\right)^{2}=\dfrac{8\pi G_{N}}{3}{\rho_{\rm vac}}.$ (34)
In the expression (3) for $\rho_{\rm vac}$}, by dimensional analysis $\mathfrak{g}$ is a dimensionless coupling constant. We assume $m_{\rm z}$ is the physical renormalized mass and ignore its potential RG flow. There are good reasons to expect that $\mathfrak{g}$ has an RG flow, based on the comments in ALrhovac1 ; ALrhovac2 ; LeClairUniverse and equation (6) where $\alpha=e^{2}/4\pi\hbar c$ is the fine structure constant with a known RG flow coming from QED. If we associate $\mathfrak{g}$ with a marginal coupling, $\mathfrak{g}$ could be marginally relevant or irrelevant. Let us assume the following beta function, where $\mu$ is an energy scale, and increasing $\mu$ corresponds to an RG flow to higher energies [^15] The factor of $1/2\pi$ is motivated by standard definitions of fine-structure constants $\alpha$, such as $e^{2}/(4\pi\hbar c)$ for QED. If $\mathfrak{g}=\alpha=\dfrac{g^{2}}{4\pi\hbar c}$ where $g$ is a gauge coupling with beta-function $\beta_{g}=\dfrac{dg}{d\log\mu}=b\dfrac{g^{3}}{16\pi^{2}}$, then $\beta_{\mathfrak{g}}=b\mathfrak{g}^{2}/2\pi$. For pure QED with $N_{f}$ species of Dirac fermions with $g=e$, $\mathfrak{g}_{0}\approx 1/137$ and $b=4N_{f}/3$. For $\lambda\phi^{4}$ theory, if $\mathfrak{g}=\lambda$ then $b=9/8\pi$. See Peskin .
$\mu\partial_{\mu}\mathfrak{g}=\dfrac{b}{2\pi}\mathfrak{g}^{2}.$ (35)
The above beta function is typically just the 1-loop approximation, however we will argue below, based on cosmological data, that the coupling $\mathfrak{g}$ is presently small so that a 1-loop approximation is justified. We assume $\mathfrak{g}>0$ for a positive $\rho_{\rm vac}$} so that we are dealing with de Sitter space. For $\mathfrak{g}>0$, if $b$ is positive the marginal coupling $\mathfrak{g}$ increases at higher energies, namely it is marginally irrelevent, whereas if $b$ is negative $\mathfrak{g}$ flows to zero in the UV, i.e. is asymptotically free. Let $\mathfrak{g}_{0}=\mathfrak{g}(\mu_{0})$ where $\mu_{0}$ is the cosmological energy scale today at time $t_{0}$. Integrating the beta function is straightforward:
$\dfrac{\mathfrak{g}(\mu)}{\mathfrak{g}_{0}}=\dfrac{1}{1-\widehat{b}\log(\mu/\mu_{0})},~~~~~\widehat{b}\equiv\dfrac{b\,\mathfrak{g}_{0}}{2\pi}.$ (36)
The RG parameter $\widehat{b}$ is the primary new constant for applications to cosmology. As we will see, based on cosmological data, we will propose $\widehat{b}>0$, corresponding to a marginally irrelevant coupling, and unless otherwise stated, below $\widehat{b}$ is positive. The pole at $1-\widehat{b}\log(\mu/\mu_{0})=0$ is commonly referred to as a Landau pole, and signifies the theory is not UV complete. It will play an important role below where we will return to addressing its significance.
From the expression (3) one has
$\dfrac{{\rho_{\rm vac}}(\mu)}{{\rho_{\rm vac}}(\mu_{0})}=\dfrac{\mathfrak{g}_{0}}{\mathfrak{g}(\mu)},$ (37)
since we assume $m_{\rm z}$ is already renormalized to the physical mass. Thus $\rho_{\rm vac}$} in (34) should be replaced by ${\rho_{\rm vac}}(\mu)$. In order to express this equation in terms of ${\rho_{\rm vac}}={\rho_{\rm vac}}(\mu_{0})$ today, it is meaningful to incorporate the RG flow into an induced flow for Newton’s constant:
$\left(\dfrac{\dot{a}}{a}\right)^{2}={\cal H}(t;\mu)^{2}\equiv\dfrac{8\pi{\cal G}(\mu)}{3}{\rho_{\rm vac}}(\mu_{0}),$ (38)
where
${\cal G}(\mu)={\cal G}(\mu_{0})\dfrac{\mathfrak{g}_{0}}{\mathfrak{g}(\mu)}={\cal G}(\mu_{0})\left(1-\widehat{b}\log(\mu/\mu_{0})\right).$ (39)
It is meaningful to incorporate the RG flow for $\rho_{\rm vac}$} into an induced flow for $G_{N}$ since we view $G_{N}$ as being determined in the far future where vacuum energy dominates and only afterwards incorporate adding matter and energy as we will do below. If one interprets ${\cal G}(\mu)$ in (39) as an energy scale dependent Newton’s constant, then one sees that for $\widehat{b}>0$, the strength of gravity decreases as one increases the energy scale $\mu$. At extremely low energy scales $\mu$, ${\cal G}(\mu)\to\infty$. The quantity ${\cal H}(t;\mu)$ is interpreted as the solution to Einstein’s equations with a Newton’s constant ${\cal G}(\mu)$ that depends on the energy scale $\mu$. Then based on (34) one has
$\dfrac{{\cal H}(t;\mu)}{{\cal H}(t;\mu_{0})}=\sqrt{\dfrac{{\cal G}(\mu)}{{\cal G}(\mu_{0})}}=\sqrt{1-\widehat{b}\log(\mu/\mu_{0})}.$ (40)
III.2 Tying the energy scale $\mu$ to the temperature in cosmology
The natural energy scale $\mu$ in cosmology is based on its temperature $T$. In the $\Lambda{\rm CDM}$ model this temperature is tracked by the cosmic microwave background (CMB). Thus far we have only considered vacuum energy with no radiation nor matter, and will turn to the $\Lambda{\rm CDM}$ model below. As a warm up exercise, let us assume the standard relation between temperature and the scale factor: [^16] For the $\Lambda{\rm CDM}$ model, $T_{0}=2.7K$, thus $\mu_{0}=k_{B}T_{0}=2.35\times 10^{-4}\,{\rm eV}$.
$T(t)=\dfrac{T_{0}}{a(t)},$ (41)
since as we will see below the main features will persist with the addition of matter. As is also standard, we express $a(t)$ in terms of the redshift $z$, $a(t)\equiv 1/{(1+z(t))}$, such that
$\mu/\mu_{0}=T/T_{0}=1+z(t).$ (42)
We thus identify $\mu_{0}=k_{B}T_{0}$, and $z=0$ with $\mu_{0}$. If we identify $G_{N}={\cal G}(\mu_{0})$, then
${\cal G}(\mu)=G_{N}\left(1-\widehat{b}\log(1+z)\right)=G_{N}\left(1+\widehat{b}\log(a)\right).$ (43)
Note that $G_{N}=0$ when $a=a_{\rm min}=e^{-1/\widehat{b}}$. For smaller $a(t)<a_{\rm min}$, in principle ${\cal G}(\mu)$ could change sign, which would signify a transition to anti-de Sitter space, however as we will see, $a(t)$ never drops below $a_{\rm min}$ and the logarithmic branch cut from $a=0$ is never reached.
Turning to (38), $\dot{a}/a=H$ is no longer constant since $\mu$ is time dependent according to (42). Taking the square-root we wish to solve
$\left(\dfrac{\dot{a}}{a}\right)=\pm\sqrt{1+\widehat{b}\log a(t)}~\left(\dfrac{8\pi G_{N}\rho_{{\rm vac};0}}{3}\right)^{1/2}$ (44)
where $\rho_{{\rm vac};0}$ is $\rho_{\rm vac}$} at the scale $\mu_{0}$. This additional kind of time dependence does not originate directly from Einstein’s equations but rather from the energy scale $\mu$ dependence of Newton’s constant and $\mu$ being tied to the temperature through equation (42). This leads to a significant and interesting modification of the de Sitter space solution to $a(t)$. Let us first chose the positive sign ++ on the RHS of (44) such that $\dot{a}>0$ in the future. As we will show, choosing the minus sign will be relevant for the far past. The equation (38) can be explicitly integrated, yielding a “inverted gaussian” correction to de Sitter space, which we will simply refer to as gaussian de Sitter space:
$a(t)=e^{-1/\widehat{b}}\,\exp\left[\dfrac{1}{\widehat{b}}\left(\dfrac{\widehat{b}H_{0}}{2}(t-t_{0})+1\right)^{2}\right]=\exp\left[H_{0}(t-t_{0})+\widehat{b}H_{0}^{2}(t-t_{0})^{2}/4\right].$ (45)
Above, $H_{0}$ equals $\dot{a}/a$ at the present time $t_{0}$, since the constant of integration is chosen such that $a(t_{0})=1$. One sees that near the present time, namely $t-t_{0}$ small, the second term in the exponential $\propto(t-t_{0})^{2}$ is suppressed, and if $\widehat{b}=0$ one recovers the usual de Sitter solution $a(t)=e^{H(t-t_{0})}$ with $H=H_{0}$.
The most interesting feature of the gaussian de Sitter solution (45) is that there is no singularity $a=0$ for all times. In spite of the Landau pole, the solution is still regular, since the argument of the square root, $1+\widehat{b}\log a(t)$, is always positive since $a(t)>a_{\rm min}$ for all times. This minimum value of $a(t)$ is
$a(t)>a_{\rm min}\equiv e^{-1/\widehat{b}}~~\forall t~~~~~\Longrightarrow~~z_{\rm max}=\dfrac{1}{a_{\rm min}}-1=e^{1/\widehat{b}}-1.$ (46)
The time $t_{\rm min}$ where $a(t_{\rm min})=a_{\rm min}$ is equal to
$t_{\rm min}=t_{0}-\dfrac{2}{H_{0}\widehat{b}}$. (47)
Note that as $\widehat{b}\to 0$, $t_{\rm min}\to$-$\infty$ as expected for de Sitter space. In the next section we will approximate $\widehat{b}\approx 0.02$ based on the so-call Hubble tension, which leads to a very large value for $z_{\rm max}$.
Let us now turn to the far past $t\to$-$\infty$. In ordinary de Sitter space, $\lim_{t\to-\infty}a(t)=0$, although it is known this is not a true geometric coordinate singularity. For our gaussian de Sitter space, the solution (45) is formally still valid for $\infty<t<t_{\rm min}$, and thus valid for all times. In fact
$\lim_{t\to\pm\infty}a(t)=\infty.$ (48)
This is a consequence of a stronger relation between the past and future. One can see from the solution (45) that it has the symmetry
$a(t)=a(-t+2t_{\rm min}).$ (49)
It is important to note that the above symmetry is independent of the present time $t_{0}$. At the time $t_{\rm min}$,   $a(t_{\rm min})=a_{\rm min}$ is self-dual. Recall that at the time $t_{\rm min}$ with scale factor $a_{\rm min}$, the effective Newton constant ${\cal G}(\mu)$ vanishes such that the solutions of (44) with ++ verses - agree, thus for $t<t_{\rm min}$ one should take the solution with the minus sign, where $a(t)$ decreases.
To summarize, the solution (45) is valid for all times $\infty<t<\infty$ where $a(t)>a_{\rm min}$. The complete history of such a Universe is that in the far past $t\to$-$\infty$ the scale factor $a(t)$ is infinitely large and starts to compress down to $a_{\rm min}$ which occurs at a time $t_{\rm min}$. Beyond this time, $t>t_{\rm min}$, the Universe expands until the scale factor $a(t)$ is again infinite at $t=+\infty$. This takes an infinite amount of time, and avoids any geometric singularities at $a=0$. There is no “Big Bang” that occurs at time $t_{\rm min}$, rather at this time the Universe smoothly transitions from a compression to an expansion, and $t_{\rm min}$ is just the time when the Universe is hottest. The Big Bang is now better described as a Swing. An picturesque analogy is the harmonic pendulum with the bottom of the swing corresponding to $a_{\rm min}$, and at the present time the Universe is on the up swing. Based on (49), a consistent possibility is that this process then repeats itself, such that Universe is in a sense oscillating about its vacuum like a pendulum harmonic oscillator. At the top of the swing the Universe is at its largest. If this is the case, then the age of this gaussian de Sitter Universe is even more infinite.
IV Adding matter: the $\Lambda{\rm CDM}$ model re-examined and the Hubble Tension
Thus far we have only considered vacuum energy density $\rho_{\rm vac}$}. We gave a heuristic derivation of Einstein’s equations in this context from the gravitational Casimir effect we formulated above. The Universe also contains radiation and matter, which we view as excitations above the vacuum, which must be included in $T_{\mu\nu}$. Local conservation requires vanishing of the covariant divergence $\nabla^{\mu}T_{\mu\nu}=0$, which is automatic due to the Bianchi identities. Thus, for instance, Newton’s universal law for the gravitational force between two masses is subsumed as a consequence. At this point one must deal with the usual Einstein’s equations except with an energy dependent Newton’s constant ${\cal G}(\mu)$:
$\left(\dfrac{\dot{a}}{a}\right)^{2}=\dfrac{8\pi}{3}{\cal G}(\mu)\,\rho_{\rm total},$ (50)
with ${\cal G}(\mu)$ defined in (43). For the moment, let us not tie the energy scale to $\mu$ to $z(t)$ as we did in (42). It is conventional to express (50) as follows
$H^{2}\equiv\left(\dfrac{\dot{a}}{a}\right)^{2}=H_{0}^{2}\left(1+\widehat{b}\log a\right)\left(\dfrac{\Omega_{\rm rad}}{a^{4}}+\dfrac{\Omega_{\rm m}}{a^{3}}+\Omega_{\Lambda}\right),$ (51)
where by definition $a(t_{0})=1$ at the present time $t_{0}$, $H(t_{0})=H_{0}$, and $\Omega_{\Lambda}$ is the $\rho_{\rm vac}$} contribution. The zero curvature ($k=0)$ motivated in Section II implies $\Omega_{\rm rad}+\Omega_{\rm m}+\Omega_{\Lambda}=1$. In the $\Lambda{\rm CDM}$ model one ignores radiation, i.e. $\Omega_{\rm rad}\approx 0$, as this is a well-justified approximation during this epoch.
There is a deepening discrepancy in the $\Lambda{\rm CDM}$ model based on relatively recent astrophysical measurements referred to as the “Hubble tension”. The ideas in this article offer a potential resolution which we now present. As usual, let $H_{0}=\dot{a}/a$ be the Hubble constant today. It should be a fixed constant regardless of the manner in which it is measured. The discrepancy arises from two very different kinds of measurements. The first comes from “local” measurements based on Type Ia supernovae, where typical redshifts are relatively low, in the range z=0.020.15z=0.02–0.15 (SHOES SHOES .). The other determination of $H_{0}$ is based on Cosmic Microwave Background data at much higher redshift $z=1100$. (Planck Planck ). The two contradictory values reported are
$H_{0;{\rm CMB}}=67.4\pm 0.5~{\rm km/s/Mpc},~~~~~H_{0;{\rm SN}}=73.0\pm 1.0~{\rm km/s/Mpc},$ (52)
which differ significantly enough for some prominent researchers to question the $\Lambda{\rm CDM}$ model, with suggestions that this could signify beyond standard model physics HubbleTensionRef . In the latter it was suggested that this could be due to a variable Dark Energy component, but no specific theoretical model was proposed.
Our interpretation of the discrepancy in these two values for $H_{0}$ in (52) is the following. Consider an observer making measurements of $H(t)$ at an earlier time, where their measurements are in a small range of higher redshift $z$ than today where today $z=0$. At this epoch, ${\cal G}(\mu)$ is effectively lower according to (43). If this observer fits their data to (51) they will predict a lower value of $H_{0}$ compared to a fit based on data taken at $z\approx 0$. The CMB measurements probe the state of the Universe at this earlier time, thus though CMB measurements are made at the present time $t_{0}$, they reflect a hypothetical observer making measurements at an earlier epoch where $z\approx 1100$. We thus propose based on (40)
$\dfrac{H_{0;{\rm CMB}}}{H_{0;{\rm SN}}}\approx\sqrt{1-\widehat{b}\log(1+z)},~~~~{\rm for}~z=z_{\rm CMB}=1100,$ (53)
where we have set $z\approx 0$ for the supernovae measurements. Based on the measured values in (52), the above leads to $\widehat{b}\approx 0.02$.
Let now turn to explicitly solving (51) in “real time”, namely with ${\cal G}(\mu(t))$ given in (42), as we did for pure vacuum energy with the result (45). Before even integrating the equation (50), one can immediately see that there are no real solutions unless $1+\widehat{b}\log a>0$. Thus there is still a minimal scale factor $a(t)>a_{\rm min}$ which is the same as in (46) and attributed to the Landau pole. This is a robust conclusion, regardless of whether $\rho_{\rm rad}$ is incorporated as the dominant form of energy in the very early Universe. The above value for $\widehat{b}$, leads to a very high value for the maximal red shift: [^17] This is not an unreasonable value for $\widehat{b}$, in that $\mathfrak{g}_{0}=2\pi\widehat{b}/b=0.12/b$, and typically $b={\cal O}(1)$. For comparison, for QED $b=4N_{f}/3$ for $N_{f}$ species of Dirac fermions, and if $\mathfrak{g}_{0}=1/137$ then $\widehat{b}=.002N_{f}$ which is about 10 times smaller than in (54) and $z_{\rm max}\approx 10^{500}$ is far beyond the Planck scale.
$\widehat{b}\approx 0.02~~~\Longrightarrow~~a_{\rm min}\approx 2\times 10^{-22},~~~~z_{\rm max}\approx 5\times 10^{21}.$ (54)
This is far above what has been directly probed by the CMB, the latter being $z_{\rm CMB}=1100$. Thus our proposed RG flow for Newton’s constant implies only small logarithmic corrections to the $\Lambda{\rm CDM}$ model. For instance it doesn’t significantly alter what was previously considered as the age of the Universe, as we will show below.
To further justify the last statement, let us turn to the $\Lambda{\rm CDM}$ model where radiation $\Omega_{\rm rad}=0$ as a good approximation,. In order to probe potential singularities at early times where the matter dominates due to the $1/a^{3}$ in (50), let us neglect the $\Omega_{\Lambda}$ term as an approximation. Then one wishes to integrate the equation
$\left(\dfrac{\dot{a}}{a}\right)^{2}=H_{0}^{2}\,\left(1+\widehat{b}\log a\right)\,\dfrac{\Omega_{m}}{a^{3}}.$ (55)
When $\widehat{b}=0$, this is easily integrated:
$a(t)^{3/2}=\tfrac{3}{2}H_{0}\sqrt{\Omega_{m}}\left(t-t_{0}+\dfrac{2}{3H_{0}\sqrt{\Omega_{m}}}\right),$ (56)
where $a(t_{0})=1$ is an initial condition. There exists a time $t_{\rm min}$ where aa=0: [^18] In the cosmology literature the convention is to shift $t\to t+t_{\rm min}$ such that $a=0$ at the shifted time $t=0$. The difference $t_{0}-t_{\rm min}$ essentially represents the age of the Universe if one assume the Universe came into existence at time $t_{\rm min}$, which is questioned in this article.
$a(t_{\rm min})=0~~{\rm for}~~t_{\rm min}=t_{0}-\dfrac{2}{3H_{0}\sqrt{\Omega_{m}}}.$ (57)
For $\widehat{b}\neq 0$ the equation (55) can also be explicitly integrated. As for the pure vacuum energy case of the last section, taking the square-root of (55) introduces a $\pm$ sign as in (44), where +/+/- corresponds to the far future/past. For the ++ sign the solution $a(t)$ can be expressed implicitly in terms of the imaginary error function ${\rm erfi}(x)\equiv{\rm erf}(ix)/i$:
${\rm erfi}\sqrt{\dfrac{3}{2\widehat{b}}\left(1+\widehat{b}\log a(t)\right)}=(t-t_{0})H_{0}\,\,\sqrt{\dfrac{3\widehat{b}\,\Omega_{m}}{2\pi}}\,e^{3/2\widehat{b}}+{\rm erfi}\sqrt{\dfrac{3}{2\widehat{b}}}$, (58)
where again we have imposed $a(t_{0})=1$. The argument of the $\rm erfi$} function on the LHS must be real otherwise the $\rm erfi$} function is imaginary. Thus due to the square-root branch cut $\sqrt{1+\widehat{b}\log a}$ inside $\rm erfi$}, there is a minimal value of $a(t)$ again given by (46). This minimal value of $a(t)$ occurs at a time $t_{\rm min}$ where $a(t_{\rm min})=a_{\rm min}$, which is determined by when the LHS of the above equation is zero since ${\rm erfi}(0)=0$:
$a(t_{\rm min})=a_{\rm min}=e^{-1/\widehat{b}}$ (59)
where
$t_{0}-t_{\rm min}=\dfrac{1}{H_{0}}e^{-3/2\widehat{b}}\sqrt{\dfrac{2\pi}{3\widehat{b}\,\Omega_{m}}}{\rm erfi}\sqrt{\dfrac{3}{2\widehat{b}}}=\dfrac{2}{3H_{0}\sqrt{\Omega_{m}}}\left(1+\tfrac{1}{3}\,\widehat{b}+\tfrac{1}{3}\,\widehat{b}^{2}+\tfrac{5}{9}\,\widehat{b}^{3}+\ldots\right).$ (60)
The solution (58) is again valid for all times $\infty<t<\infty$ due to a symmetry between the far past and far future. Using ${\rm erfi}(-x)=-{\rm erfi}(x)$, one can show that equation (49), which equates $a(t)$ with $a(-t+2t_{\rm min})$, remains valid with $t_{\rm min}$ given in (60), as does (48).
Based on the above analysis of our scenario, let us comment on the possible implications for the very early Universe. Henceforth, by “early Universe” we mean around the time $t_{\rm min}$, where the expansion phase begins. Recall in our model, time $t$ actually extends to $t=$-$\infty$, and this is the earliest Universe. The salient features of the evolution of $a(t)$ from $t=$-$\infty$ to $t=+\infty$ is essentially the same as for the pure vacuum energy case, described in the last paragraph of the last section. Namely at $t=$-$\infty$, $a(t)=\infty$ and starts to decrease until it reaches $a_{\rm min}$ at time $t_{\rm min}$, then begins to increase back to $\infty$ as $t\to+\infty$. At no time in this evolution is $a=0$, again avoiding geometric curvature singularities. There is no singular Big Bang at any time, rather at time $t_{\rm min}$ the universe is hottest since the redshift $z$ is at its maximum. The quantity $t_{0}-t_{\rm min}$ does not represent the age of the Universe, but rather the time elapsed since the scale factor was at its minimum. The age of the Universe is actually infinite, which we find appealing, since the Universe does not originate from an incomprehensible singularity at $a=0$. As described above for the case of only Vacuum energy, the time evolution of the Universe is better described as a back and forth swing of a pendulum, where the bottom of the swing occurs at $a(t_{\rm min})=a_{\rm min}$, and we are currently on an upswing.
The details of this evolution are very sensitive to the value of $\widehat{b}$ due to the exponentials in (54). For purposes of illustration, let us assume $\widehat{b}=0.02$ as inferred from the Hubble tension. First, one sees from (60) that for small $\widehat{b}$ this amounts to a small corrections to $t_{0}-t_{\rm min}$ which is normally associated to the age of the Universe. If our estimate of $\widehat{b}$ in (54) is approximately correct, then this value for $z_{\rm max}$ is deeply into the radiation dominated era, since the value of $z=z_{\rm eq}$ where matter and radiation are equivalent in density is only about $z_{\rm eq}\approx 3400$. At times much earlier than during the $\Lambda{\rm CDM}$ era, equation (42) is not necessarily valid, however if we simply assume it extends to the Planck scale with $T_{0}=2.7$K, then $z_{\rm planck}\approx 5\times 10^{31}$, which is much higher than $z_{\rm max}$ by 10 orders of magnitude. This indicates that the Planck scale plays no role in determining $z_{\rm max}$, which is to say quantum gravity does not play a role.
For our perspective on gravity, earlier times $t<t_{\rm min}$ where $a(t)<a_{\rm min}$ do not exist in our model. In the standard cosmology this earlier era is thought to be described by inflation. Inflation is postulated as a way to get from the singularity at $a=0$ to a hot Big Bang, however it is not yet universally accepted as being a complete theory. Very large numbers are naturally involved since the singularity at $a=0$ formally corresponds to $z=\infty$. Typical models require inflation to last for 50-60 e-foldings to inflate the Universe to the proper initial conditions from $a\approx 0$, then transition to a re-heating scale to bring up the temperature to that of a hot Big Bang, and this rapid expansion makes the Universe flat and homogeneous. The range of redshift $z$ over which inflation hypothetically occurs is strongly model dependent, especially on what the re-heating temperature is, and the duration of inflation, marking its beginning and end, can range from $10^{22}<z_{\rm inflation}<10^{53}$, where the upper limit can even be significantly larger than $z_{\rm Planck}$. [^19] During hypothetical inflation, the redshift $z$ is not given by the formula (42), but based on the definition of the scale factor $a=1/(1+z)$. It is noteworthy that at the lower end of this range based on lower reheating temperatures, namely where inflation ends at $z_{\rm inflation}=10^{22}$, is close to our $z_{\rm max}=5\times 10^{21}$. Since this inflationary epoch does not actually exist in our model, one should question whether inflation is necessary in the standard model of cosmology. At $t=t_{\rm min}$ the Universe is already expanded since $a(t)>a_{\rm min}$, and the Universe is already flat with $k=0$. In Section II we explained why the spatial curvature $k$ is naturally zero since we view Gravity as arising from vacuum energy in flat Minkowski space. The value of $a_{\rm min}=e^{-1/\widehat{b}}$ perhaps plays a role equivalent to hypothetical inflation models since it represents roughly 50 e-foldings since $e^{1/\widehat{b}}\approx e^{50}$ for our estimate of $\widehat{b}\approx 0.02$.
V Closing Remarks
Having already summarized the unconventional perspective on Gravity developed in this article and its implications in the Abstract and Introduction, we close with some suggestions for further investigations and discuss some open questions. There are currently many open avenues for testing (and falsifying) our main proposals.
$\bullet$   Above we considered implications for the cosmology based on the $\Lambda{\rm CDM}$ model. The energy scales involved are relatively low since $\mu_{0}=k_{B}T_{0}=2.35\times 10^{-4}\,{\rm eV}$ for $T_{0}=2.7$ K. If our proposed energy dependent Newton’s constant ${\cal G}(\mu)$ in equation (43) is correct, then it is in principle possible to confirm it in some bench-top Cavendish like experiments as a function of temperature. Suppose such experiments are carried out in a small range of temperature $\Delta T$ about a fixed temperature $\widetilde{T}$. From (39)
${\cal G}(T)={\cal G}(\widetilde{T})\left(1-\widehat{b}\log(T/\widetilde{T})\right)\approx{\cal G}(\widetilde{T})\left(1-\dfrac{\widehat{b}}{\widetilde{T}}\,\Delta T\right).$ (61)
As discussed above, the effective Newton’s constant decreases with increased temperature. If $\widetilde{T}=300$ K is room temperature, then $\widehat{b}/\widetilde{T}\approx 6\times 10^{-5}$ per degree change $\Delta T$ if $\widehat{b}=0.02$, where this value of $\widehat{b}$ was inferred from the Hubble tension in (53). There are relatively few experiments along these lines, going back as far as 1923 Shaw ; Dmitriev ; Liangzao , wherein weight is measured as a function of temperature for various metals around room temperature. These experiments indeed consistently measure a decrease in the gravitational force as the temperature is increased, as our model predicts. Based on the ideas in this article, this decrease should be independent of the material, in this case a metal, since the Equivalence Principle is built into the formalism, whereas experiments find slightly different relative changes to $G_{N}$ depending on the metal. This could be due to experimental uncertainties as a result of buoyancy, thermal expansion, and other factors which are difficult to take into account. In any case a linear fit of this experimental data in $\Delta T$ according to (61) yield variations of about $10^{-5}$ per degree, which is roughly consistent with our estimate of $\widehat{b}/\widetilde{T}$ for $\widehat{b}\approx 0.02$ and $\widetilde{T}=300$K. [^20] The article Liangzao does not attempt a fit to their data. We made a crude fit of their tabular results to the functional form (61) and found $\widehat{b}/\widetilde{T}$ to be $0.8\times 10^{-5}$ for gold and $1.4\times 10^{-5}$ for nickel, however a closer inspection and interpretation of this data beyond our crude fit is warranted. At lower reference temperatures $\widetilde{T}$, the effect is larger and this motivates carrying out such Cavendish type experiments at lower temperatures.
$\bullet$   The induced RG flow for Newton’s constant that we proposed as a consequence of the RG flow of the coupling $\mathfrak{g}$ in $\rho_{\rm vac}$} in (3) led to an avoidance of a geometric curvature singularity at a scale factor $a=0$. There are various theorems asserting that such singularities are unavoidable in cosmology and black holes Penrose ; HawkingPenrose . The latter theorems assume a constant Newton’s constant $G_{N}$, and should be revisited to account for an effective energy scale dependent $G_{N}$ as proposed in this article. This could potentially lead to a resolution of the singularity at the origin of the Schwarzschild black hole solution, in a way analogous to the minimal scale factor $a_{\rm min}$ above.
$\bullet$   If the varying $G_{N}$ proposed in this article is correct, it is testable by other kinds of measurements. Let us speculate on some possible directions or topics for further study. We already mentioned above that in principle this varying $G_{N}$ can be measured in bench-top Cavendish experiments as a function of temperature, and that the very few existing experiments carried out at room temperature are roughly consistent with the parameters we inferred from the Hubble tension. For cosmology, our formula (39) for the scale dependent Newton’s constant ${\cal G}(\mu)$ implies that the effective Newton’s constant is effectively time dependent if one ties $\mu$ to the time-dependent scale factor as in (42). There are known experimental bounds on the time variation of $G_{N}$, usually reported as $\dot{G}_{N}/G_{N}$. Based on (43) one has
$\dfrac{\dot{G}_{N}}{G_{N}}=H\,\widehat{b}$, (62)
where as above $H=\dot{a}/a$. The strongest experimental constraints come from pulsar timing Taylor which constrains $\dot{G}_{N}/G_{N}<2\times 10^{-12}/{\rm yr}$. LIGO provides weaker constraints $10^{-9}/{\rm yr}$ LIGO . These pulsars exist at relatively low redshift $z$. Thus taking $H=H_{0}$ in the above equation (62), one finds $\dot{G}_{N}/G_{N}\approx 7\times 10^{-11}\times\widehat{b}\,/{\rm yr}$. For $\widehat{b}\approx 0.02$, as we proposed based on Hubble tension data, $\dot{G}_{N}/G_{N}\approx 1.4\times 10^{-12}\,/{\rm yr}$ which is very close to the bound based on pulsars. There are many other testable predictions in cosmology in addition to the Hubble tension and those mentioned above. A varying $G_{N}$ may shift CMB acoustic peaks due to modified expansion at $z\sim 1100$, which is testable with Planck or the Simons Observatory data. The minimal scale factor $a_{\rm min}\sim 10^{-22}$ in eq. (46) could produce a primordial gravitational wave spectrum detectable by LIGO or eventually the space based interferometer LISA. A weaker $G_{N}$ at higher redshift $z$ could delay stellar collapse, favoring primordial black holes. In fact, based on (32), for a weaker $G_{N}$, black holes are smaller for a given mass. Very recently, supermassive black holes, with mass on the order of 40 million solar masses have been discovered using the James Webb Space Telescope JWST which were formed at an earlier time than expected, and this may be an interesting topic for further study in the context of a varying $G_{N}$.
$\bullet$   As stated in the Introduction, although the formula (3) is well-motivated by the work ALrhovac1 ; ALrhovac2 , the QFT underlying this formula for $\rho_{\rm vac}$} has not been precisely identified, in particular the zeron with mass $m_{\rm z}$ has not yet been identified as a known particle, though Majorana fermions are viable candidates. This underlying QFT is incomplete in the UV since it relies on the RG beta function (35) for a marginally irrelevant coupling, where the incompleteness is manifested as a Landau pole at very high energies. However we showed that the time evolution of the scale factor $a(t)$ smoothly traverses this pole through a swing rather than a bang. The UV completeness issue here is certainly more tractable than the UV incompleteness of quantum gravity, since it is a UV issue of QFT in flat Minkowski space where such problems are better understood. Landau poles are commonplace in elementary particle physics and condensed matter physics, since they are generic to marginally irrelevant perturbations. See for instance LandauPole1 ; LandauPole2 ; LandauPole3 and references therein. Although QED itself is not relevant to this article, let us mention that it also is not UV complete due to a Landau pole, however in this context it occurs at much higher energies than our proposed $z_{\rm max}$, and is thus commonly viewed as a purely academic issue, and is furthermore complicated by chiral symmetry breaking. (See footnote 12).